# Understanding Catastrophic Forgetting in Language Models via Implicit Inference

### v1

@@ Abstract
Fine-tuning (via methods such as instruction-tuning or reinforcement learning from human feedback) is a crucial step in training language models to robustly carry out tasks of interest. However, we lack a systematic understanding of the effects of fine-tuning, particularly on tasks outside the narrow fine-tuning distribution. In a simplified scenario, we demonstrate that improving performance on tasks within the fine-tuning data distribution comes at the expense of suppressing model capabilities on other tasks. This degradation is especially pronounced for tasks “closest” to the fine-tuning distribution. We hypothesize that language models implicitly infer the task of the prompt corresponds, and the fine-tuning process predominantly skews this task inference towards tasks in the fine-tuning distribution. To test this hypothesis, we propose Conjugate Prompting to see if we can recover pretrained capabilities. Conjugate prompting artificially makes the task look farther from the fine-tuning distribution while requiring the same capability. We find that conjugate prompting systematically recovers some of the pretraining capabilities on our synthetic setup. We then apply conjugate prompting to real-world LLMs using the observation that fine-tuning distributions are typically heavily skewed towards English. We find that simply translating the prompts to different languages can cause the fine-tuned models to respond like their pretrained counterparts instead. This allows us to recover the in-context learning abilities lost via instruction tuning, and more concerningly, to recover harmful content generation suppressed by safety fine-tuning in chatbots like ChatGPT.
footnotetext: Code available at https://github.com/kothasuhas/understanding-forgetting
@@ 1 Introduction
The development of large language models (LLMs) typically involves two stages—pretraining (next token prediction) on vast text corpora and fine-tuning on carefully curated datasets to adapt the pretrained model to the application of interest. The fine-tuning stage is critical in enabling language models to output helpful text, and there has been significant interest in various fine-tuning methods such as instruction-tuning  (Ouyang et al., 2022; Wei et al., 2022; Mishra et al., 2022; Chung et al., 2022; Taori et al., 2023) and reinforcement learning from human feedback (Christiano et al., 2023; Stiennon et al., 2022; Bai et al., 2022; Ziegler et al., 2020).
@@
One fundamental concern is that fine-tuning datasets are considerably smaller and less diverse than web-scale pretraining datasets (Raffel et al., 2020; Arivazhagan et al., 2019; Gao et al., 2021), and there is always a risk that the fine-tuned model “catastrophically forgets” (McCloskey & Cohen, 1989) how to solve problems that the pretrained model could solve. Such a gap has been reported as an “alignment tax” in works such as Ouyang et al. (2022) and Bai et al. (2022), but there is no clear understanding of what these trade-offs are and how to mitigate them. Given the importance of the fine-tuning process, it is imperative to build a systematic understanding of the effects.
@@
In this work, we introduce a synthetic setup to understand the effects of fine-tuning. Building on the prior work of in-context learning linear functions (Garg et al., 2023) by pretraining transformers (Vaswani et al., 2017) on a large number of weight vectors, we show that the resulting transformers can be sub-optimal when evaluated on a few specific weight vectors of special interest. This mirrors real-world settings where the pretraining data is uncurated data from the web, while there are some “natural” tasks of special interest, like question answering. We find that fine-tuning on the weights (tasks) of interest enables transformers to achieve optimal performance on these tasks but comes at the cost of performing worse on other tasks (Figure 4).
@@
We further examine the performance degradation induced by fine-tuning and observe a striking structure in terms of which tasks are adversely affected. The most affected tasks are outside but still “close” to the fine-tuning distribution as measured by their likelihood under the fine-tuning distribution (Figure 5). In other words, the fine-tuned model performs more like the pretrained model on tasks that are far from the fine-tuning distribution. We hypothesize this is because transformers are implicitly “inferring” the task before solving the corresponding task. The fine-tuning process might not significantly change the capability to solve tasks (inside and outside the fine-tuning distribution), but rather skew the task inference heavily towards the fine-tuning task distribution, disproportionately affecting tasks close to but outside the fine-tuning distribution.
@@
Assuming this framework, we can recover the suppressed pretraining capability by forcing a prompt to look very different from the fine-tuning distribution, helping the fine-tuned model infer that the task is outside the fine-tuning distribution. To do this, we propose Conjugate Prompting. For a prompt $P$ outside the fine-tuning distribution, we prompt the language model with prompt $P^{\prime}$ such that (i) $P^{\prime}$ is less likely under the fine-tuning distribution and (ii) the solution to prompt $P$ can be easily recovered from the solution to prompt $P^{\prime}$. Since $P^{\prime}$ is farther from the fine-tuning distribution than $P$, the fine-tuned model will solve $P^{\prime}$ with the pretrained capability, allowing us to extract a better solution for the original task $P$. We test our conjugate prompting method in the linear regression setup and find that it alleviates some of the trade-offs induced by fine-tuning (Figure 6).
@@
Drawing inspiration from the synthetic experiments, we validate whether fine-tuning affects real language models in the same manner. Since fine-tuning datasets are curated and primarily in English, we apply conjugate prompting with language translation to lower the likelihood of being drawn from the fine-tuning distribution while preserving the core task. We construct a problem where it is ambiguous whether the correct task is in-context learning or following an instruction and find that instruction-tuning suppresses in-context learning. Across 5 models and 4 non-English languages (with 2 additional transformations), conjugate prompting recovers the pretrained capability of in-context learning. We also consider the problem of harmful content generation where language models like ChatGPT are actively fine-tuned to suppress the capability of answering harmful instructions: here it is in an adversary’s interest to recover suppressed capabilities of pretraining. We find that conjugate prompting can still circumvent the fine-tuned capability of refusal of harmful instructions and recover the pretrained capability of following the instruction.
@@
Figure 1: How does fine-tuning affect language models? When pretrained over data that contains the orange task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$ and the blue task $\color[rgb]{0.265625,0.4453125,0.76953125}T_{2}$, a model may infer a prompt $P$ is from task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$ and solve the task accordingly. When fined-tuned over task $\color[rgb]{0.265625,0.4453125,0.76953125}T_{2}$, the model may no longer perform task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$. We hypothesize that this might not mean the task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$ is forgotten, but rather just the implicit task inference is shifted away from $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$. Leveraging this viewpoint, we provide conjugate prompting to recover pretrained model behavior by countering the change in implicit task inference, shedding light onto the nature of catastrophic forgetting.
@@ 2 Linear Regression Experiments
We explore a simple synthetic setup where we train transformers to in-context learn linear functions. Our setup tries to mirror the structure of large language model training by pretraining over a broad class of many tasks, from the distribution $\mathcal{D}_{\text{cont}}$, and a special set of few tasks, from the distribution $\mathcal{D}_{\text{disc}}$ (Section 2.4). When we fine-tune the model to improve performance over $\mathcal{D}_{\text{disc}}$, the model seems to “forget” the capability to solve tasks from $\mathcal{D}_{\text{cont}}$ (Section 2.5). However, we hypothesize that these capabilities are rather “suppressed” (Sections 2.6 and 2.7), and we find that we can recover them through our principled conjugate prompting strategy (Section 2.8).
2.1 Setup: in-context learning of linear functions
@@
We consider auto-regressive models $T_{\theta}$ that take in a sequence of tokens, each in $\mathbb{R}^{d}$, and produce a real-valued output. For samples $S$ generated under $w$ as in Equation 1, we feed the model $T_{\theta}$ the sequence $\left[x_{1},y_{1},\ldots,x_{k},y_{k},x_{\text{query}}\right]$ [^1] Every $1$-dimensional token is right-padded with $d-1$ zeroes (which we will refer to as the prompt) and take its output $\hat{y}$ as a prediction of $y_{\text{query}}$. When appropriate, we will refer to the $x_{i}$’s in the prompt as $X\in\mathbb{R}^{k\times d}$ and the $y_{i}$’s as $y\in\mathbb{R}^{k}$. We train and evaluate $T_{\theta}$ with respect to a weight distribution $\mathcal{D}$ via the quadratic loss
@@
by sampling a fresh batch of $x,w,\epsilon$ in each step. Under the quadratic loss, the optimal output is $\mathbb{E}\left[f_{w}(x_{\text{query}})+\epsilon\mid X,y\right]=\langle\mathbb{E}\left[w\mid X,y\right],x_{\text{query}}\rangle$. For our model, we use a 22.4 million paramater GPT-2 style transformer. For a comprehensive explanation of our experimental setup, refer to Appendix B.5.
@@ 2.2 Gaussian prior over weights ($\mathcal{D}_{\text{cont}}$)
As noted in prior work, for most values of $\tau,\sigma$, a converged transformer’s predictions closely match the Bayes optimal predictor when evaluated on new weight vectors from the same Gaussian prior. We replicate this for $\tau=1$ in Figure 2, left.
@@ 2.3 Discrete prior over fixed weights ($\mathcal{D}_{\text{disc}}$)
The Gaussian prior spreads probability mass over a large region of weight vectors, but in real world distributions, there isn’t such a “uniform” prior over the task space. Rather, there are a few common tasks (e.g. summarization or sentiment analysis) which frequently appear in the task distribution, and pretrained LLMs rely on these priors for producing outputs (Min et al., 2022b; Wei et al., 2023b; Pan et al., 2023).
We take this scenario to the extreme and consider training over a “fixed” set of weights with the distribution $\mathcal{D}_{\text{disc}}$ sampling $w$ uniformly from $\{w_{1},\ldots,w_{n}\}$. We refer to this as the “discrete distribution”. For our experiments, we set $n=64$ and fix each $w_{i}$ as an independent sample of $\mathcal{N}(0,I_{d})$. With this new prior over the weights, ridge regression is no longer the optimal solution. The Bayes optimal estimator for $\mathcal{D}_{\text{disc}}$ is:
@@
where $\varphi\left(\cdot\right)$ is the density of the standard multivariate normal distribution (derivation in Appendix A.1). We refer to this estimator as discrete regression. After training for sufficiently many steps, we find that the Transformer achieves the same loss as the Bayes-optimal estimator $w^{*}_{\text{disc}}$, clearly outperforming ridge regression on the fixed set of weights (Figure 2, right).
@@ 2.4 Pretraining over the mixture ($\mathcal{D}_{\text{mix}}$)
We know that web-scale pretraining data is heavy-tailed, consisting of some important tasks seen often (similar to $\mathcal{D}_{\text{disc}}$), as well a large number of diverse tasks each seen rarely (similar to $\mathcal{D}_{\text{cont}}$). To best model this structure, we consider the mixture of the continuous and discrete distributions, which we refer to as the “mixture distribution”
@@
where $g(X,y)$ is the posterior probability that the prompt $X,y$ was sampled from $\mathcal{D}_{\text{disc}}$ (complete expression and derivation in Appendix A.1). Intuitively, the Bayes optimal predictor utilizes ridge regression to get $w^{*}_{\text{cont}}$ and discrete regression to get $w^{*}_{\text{disc}}$ which it appropriately weights by the posterior. We refer to this solution as mixture regression.
@@
For a given $\alpha$, we measure model performance by evaluating its loss on the continuous and discrete distributions. For mixture regression, there is a natural trade-off between performance on these distributions based on the prior determined by $\alpha$ as pictured by the black curve in Figure 3. For $\alpha$ close to 00, mixture regression heavily weights ridge regression, while for $\alpha$ close to $1$, it weights discrete regression. For intermediate $\alpha$, mixture regression can utilize the posterior to infer the distribution and get low loss on both $\mathcal{D}_{\text{cont}}$ and $\mathcal{D}_{\text{disc}}$.
Figure 3: Trade-off over training. We measure the loss over $\mathcal{D}_{\text{cont}}$ and $\mathcal{D}_{\text{disc}}$ for different models over different values of $\alpha$. The black curve, mixture regression, faces a natural trade-off over different values of $\alpha$. We also pretrain models for $\alpha\in\{0.2,0.5,0.8\}$ and measure their losses at $1000,2000,3000,4000$ and $5000$ steps. The solid red lines are trajectories over time for a fixed $\alpha$ and the dotted red lines are the trade-off for a fixed time step. Over the course of training, models approach mixture regression.
@@
We can similarly measure the loss of transformers evaluated on the continuous and discrete distributions. We find that as we train models for longer on the mixture distribution, they approach the Bayes-optimal solution of mixture regression for the respective $\alpha$. However, this convergence is very slow, especially for smaller values like $\alpha=0.2$. Moreover, even a converged model is bounded in how well it can perform on the discrete distribution due to the trade-off presented by mixture regression.
@@ 2.5 The effect of fine-tuning pretrained models
In practice, there is often a distributional mismatch between the tasks the model learned to solve during pretraining and the tasks of interest to an end user. For example, training for next token prediction over internet data doesn’t naturally respond to human-written instructions or avoid outputting toxic content. Additionally, pre-training on all tasks is not a data-efficient nor compute-efficient method to improve performance on a specific target application.
The most common solution to these problems is to fine-tune the pretrained model over the tasks of interest. We replicate this in our controlled setup by targeting performance on the fixed set of discrete tasks in $\mathcal{D}_{\text{disc}}$ which requires the model to perform discrete regression (Equation 7). Fine-tuning is necessary since pretraining is both inefficient and limited by the distributional mismatch.
@@
Fine-tuning the pretrained models from Section 2.4 over $\mathcal{D}_{\text{disc}}$ rapidly improves performance on $\mathcal{D}_{\text{disc}}$, demonstrating the utility of fine-tuning. However, this also leads to large performance drops on $\mathcal{D}_{\text{cont}}$. For example, as seen in Figure 4, when we fine-tune a transformer pretrained over $\mathcal{D}_{\text{mix}}$ for $\alpha=0.5$, the loss decreases on $\mathcal{D}_{\text{disc}}$ while it increases on $\mathcal{D}_{\text{cont}}$. This is an instance of “catastrophic forgetting” (McCloskey & Cohen, 1989), where fine-tuning a model to improve at one task causes it to worsen at other tasks.
This performance drop could imply that the model can not perform ridge regression anymore. Thanks to our synthetic setup, we investigate how fine-tuning is affecting model predictions and utilize this understanding to recover lost performance on the continuous distribution.
Figure 4: Fine-tuning improves discrete loss and hurts continuous loss. We pretrain a transformer over the mixture of $\alpha=0.2$ with $64$ discrete tasks for $5000$ steps. We fine-tune this model for $400$ steps on only $\mathcal{D}_{\text{disc}}$ as highlighted in yellow. The discrete loss rapidly decreases, while the continuous loss rapidly increases.

### v2

@@ Abstract
We lack a systematic understanding of the effects of fine-tuning (via methods such as instruction-tuning or reinforcement learning from human feedback), particularly on tasks outside the narrow fine-tuning distribution. In a simplified scenario, we demonstrate that improving performance on tasks within the fine-tuning data distribution comes at the expense of capabilities on other tasks. We hypothesize that language models implicitly infer the task of the prompt and that fine-tuning skews this inference towards tasks in the fine-tuning distribution. To test this, we propose Conjugate Prompting, which artificially makes the task look farther from the fine-tuning distribution while requiring the same capability, and we find that this recovers some of the pretraining capabilities in our synthetic setup. Since real-world fine-tuning distributions are predominantly English, we apply conjugate prompting to recover pretrained capabilities in LLMs by simply translating the prompts to different languages. This allows us to recover in-context learning abilities lost via instruction tuning, natural reasoning capability lost during code fine-tuning, and, more concerningly, harmful content generation suppressed by safety fine-tuning in chatbots like ChatGPT.
footnotetext: Code available at github.com/kothasuhas/understanding-forgetting
@@ 1 Introduction
Developing large language models (LLMs) typically involves two stages—pretraining on vast text corpora and fine-tuning on curated datasets. Fine-tuning through methods such as instruction-tuning and reinforcement learning from human feedback is critical in enabling language models to output desirable text. Since fine-tuning datasets are considerably smaller and less diverse than web-scale pretraining datasets, and there is always a risk that the fine-tuned model “catastrophically forgets” (McCloskey & Cohen, 1989) how to solve tasks that the pretrained model could solve. Such a gap has been reported as an “alignment tax” (Ouyang et al., 2022; Bai et al., 2022), but there is no clear understanding of what these trade-offs are and how to mitigate them.
@@
In this work, we introduce a synthetic setup to understand the effects of fine-tuning. Building on the prior work of in-context learning linear functions (Garg et al., 2023) by pretraining transformers (Vaswani et al., 2017) on a large number of weight vectors, we show that the resulting models can be sub-optimal when evaluated on a few weight vectors of special interest. This mirrors real-world settings where the uncurated pretraining data contains some “natural” tasks of special interest, like question answering. Fine-tuning on the weights (tasks) of interest enables transformers to achieve optimal performance on these tasks at the cost of worse performance on other tasks.
@@
We find that the most affected tasks are outside but “close” to the fine-tuning distribution as measured by their likelihood under the fine-tuning distribution. In other words, the fine-tuned model performs more like the pretrained model on tasks that are far from the fine-tuning distribution. We hypothesize this is because models are implicitly “inferring” the task before solving the corresponding task, and that fine-tuning skews model task inference more than it changes models capabilities.
@@
Assuming this framework, we can recover the suppressed pretraining capability through Conjugate Prompting. For a prompt $P$ outside the fine-tuning distribution, we prompt the language model with prompt $P^{\prime}$ such that (i) $P^{\prime}$ is less likely under the fine-tuning distribution and (ii) the solution to prompt $P$ can be easily recovered from the solution to prompt $P^{\prime}$. Since $P^{\prime}$ is farther from the fine-tuning distribution than $P$, the fine-tuned model will solve $P^{\prime}$ with the pretrained capability, allowing us to extract a better solution for the original prompt $P$. We test conjugate prompting in the linear regression setup and find that it alleviates some of the trade-offs induced by fine-tuning.
@@
Drawing inspiration from the synthetic experiments, we validate whether fine-tuning similarly affects real language models. Since fine-tuning datasets are primarily in English, we apply conjugate prompting with language translation to lower the likelihood of being drawn from the fine-tuning distribution while preserving the core task. We construct a problem that can either be solved by in-context learning or following an instruction and find that instruction-tuning suppresses in-context learning. Across 5 models and 4 non-English languages (with 2 additional transformations), conjugate prompting recovers the pretrained capability of in-context learning. We then consider a more natural form of catastrophic forgetting where fine-tuning on code leads to worse performance on a sentence entailment task testing natural language reasoning. We find that conjugate prompting improves performance on this task, sometimes even observing a slight increase in performance when prompting the fine-tuned model in non-English languages. Finally, we consider the problem of harmful content generation where chatbots like ChatGPT are fine-tuned to refuse harmful instructions: here it is in an adversary’s interest to recover suppressed pretraining capability of following the instruction. We find that conjugate prompting can circumvent the fine-tuned capability of refusal and can recover some of the pretrained capability of following the instruction.
@@
Figure 1: How does fine-tuning affect language models? When pretrained over the orange task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$ and the blue task $\color[rgb]{0.265625,0.4453125,0.76953125}T_{2}$, a model may infer a prompt $P$ is from task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$ and solve this task. When fined-tuned over task $\color[rgb]{0.265625,0.4453125,0.76953125}T_{2}$, the model may no longer perform task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$. We hypothesize that this might not mean the task $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$ is forgotten, but rather that the implicit task inference is shifted away from $\color[rgb]{0.9296875,0.48828125,0.19140625}T_{1}$. Leveraging this viewpoint, we provide conjugate prompting to recover pretrained model behavior by countering the change in implicit task inference, shedding light onto the nature of forgetting.
@@ 2 Linear Regression Experiments
We explore a synthetic setup where we train transformers to in-context learn linear functions. Our setup mirrors language model training by pretraining over a broad class of tasks from the distribution $\mathcal{D}_{\text{cont}}$ and a special set of few tasks from the distribution $\mathcal{D}_{\text{disc}}$ (Section 2.4). When we fine-tune to improve performance on $\mathcal{D}_{\text{disc}}$, the model seems to “forget” the capability to solve tasks from $\mathcal{D}_{\text{cont}}$ (Section 2.5). However, we hypothesize that these capabilities are actually “suppressed” (Sections 2.6 and 2.7) and find that we can recover them through conjugate prompting (Section 2.8).
2.1 Setup: in-context learning for linear functions
@@
We consider auto-regressive models $T_{\theta}$ that take in a sequence of tokens, each in $\mathbb{R}^{d}$, to produce a real-valued output. For samples $S$ generated under $w$ as in Equation 1, we feed $T_{\theta}$ the prompt $\left[x_{1},y_{1},\ldots,x_{k},y_{k},x_{\text{query}}\right]$ [^1] Every $1$-dimensional token is right-padded with $d-1$ zeroes and take its output as a prediction of $y_{\text{query}}$. When appropriate, we will refer to the $x_{i}$’s in the prompt as $X\in\mathbb{R}^{k\times d}$ and the $y_{i}$’s as $y\in\mathbb{R}^{k}$. We train and evaluate $T_{\theta}$ with respect to a weight distribution $\mathcal{D}$ via the quadratic loss
@@
by sampling a fresh batch of $x,w,\epsilon$ in each step. Under the quadratic loss, the optimal output is $\mathbb{E}\left[f_{w}(x_{\text{query}})+\epsilon\mid X,y\right]=\langle\mathbb{E}\left[w\mid X,y\right],x_{\text{query}}\rangle$. For our model, we use a 22.4 million paramater GPT-2 style transformer. For more experimental details, refer to Appendix C.8.
@@ 2.2 Gaussian prior over weights ($\mathcal{D}_{\text{cont}}$)
As noted in prior work ((Garg et al., 2023; Akyürek et al., 2022)), for most values of $\tau,\sigma$, a converged transformer’s predictions closely match the Bayes optimal predictor when evaluated on weight vectors from the same Gaussian prior. We replicate this for $\tau=1$ in Figure 2, left.
@@ 2.3 Discrete prior over fixed weights ($\mathcal{D}_{\text{disc}}$)
The Gaussian prior spreads probability mass over a large region of weight vectors, but in real world distributions, there isn’t such a “uniform” prior over the task space. Rather, there are a few common tasks (e.g. summarization or sentiment analysis) which frequently appear in the task distribution, and pretrained LLMs utilize these priors (Min et al., 2022b; Wei et al., 2023b; Pan et al., 2023).
We take this scenario to the extreme and consider training over a “fixed” set of weights with the distribution $\mathcal{D}_{\text{disc}}$ sampling $w$ uniformly from $\{w_{1},\ldots,w_{n}\}$. We refer to this as the “discrete distribution”. For our experiments, we set $n=64$ and fix each $w_{i}$ as an independent sample of $\mathcal{N}(0,I_{d})$. With this new prior, ridge regression is no longer optimal. The Bayes optimal estimator for $\mathcal{D}_{\text{disc}}$ is:
@@
where $\varphi\left(\cdot\right)$ is the density of the standard multivariate normal distribution (derivation in Appendix B.1). We refer to this estimator as discrete regression. After training for sufficiently many steps, we find that the Transformer achieves the same loss as the Bayes-optimal estimator $w^{*}_{\text{disc}}$, clearly outperforming ridge regression on the fixed set of weights (Figure 2, right).
@@ 2.4 Pretraining over the mixture ($\mathcal{D}_{\text{mix}}$)
We know that web-scale pretraining data is heavy-tailed, consisting of some important tasks seen often (similar to $\mathcal{D}_{\text{disc}}$), as well a large number of diverse tasks each seen rarely (similar to $\mathcal{D}_{\text{cont}}$). To best model this structure, we consider the “mixture distribution”
@@
where $g(X,y)$ is the posterior that $X,y$ was sampled from $\mathcal{D}_{\text{disc}}$ (expression in Appendix B.1). Intuitively, this predictor utilizes ridge regression to get $w^{*}_{\text{cont}}$ and discrete regression to get $w^{*}_{\text{disc}}$ which it appropriately weights by the posterior. We refer to this solution as mixture regression.
@@
We measure performance by evaluating loss on the continuous and discrete distributions, and we find a natural trade-off between performance on these distributions determined by the prior $\alpha$ (Figure 3, black curve). Mixture regression weights ridge regression for $\alpha$ close to 00 and discrete regression for $\alpha$ close to $1$. For intermediate $\alpha$, mixture regression can utilize the posterior to infer the distribution and get low loss on both $\mathcal{D}_{\text{cont}}$ and $\mathcal{D}_{\text{disc}}$ (Appendix C.5 discusses mixture regression’s ability to infer the distribution in more detail).
Figure 3: Trade-off over training. We measure the loss over $\mathcal{D}_{\text{cont}}$ and $\mathcal{D}_{\text{disc}}$ for different models over different values of $\alpha$. Mixture regression, faces a natural trade-off over different values of $\alpha$. We also pretrain models for $\alpha\in\{0.2,0.5,0.8\}$ and measure their losses at $1000$ to $5000$ steps. The solid pink lines are trajectories over time for a fixed $\alpha$ and the dotted pink lines are the trade-off for a fixed time step, showing models approaching mixture regression.
@@
As we train models on the mixture distribution, they approach the Bayes-optimal solution of mixture regression for the respective $\alpha$. However, this convergence is very slow, especially for smaller values like $\alpha=0.2$. Moreover, the trade-off bounds how well a converged model can do on the discrete distribution.
@@ 2.5 The effect of fine-tuning pretrained models
In practice, there is a distributional mismatch between the tasks learned during pretraining and the tasks of interest to an end user. For example, next token prediction over the internet doesn’t naturally respond to human-written instructions or avoid outputting toxic content. Additionally, pre-training on all tasks is not a data nor compute efficient way of improving performance on a target application.
The most common solution is to fine-tune the pretrained model over the tasks of interest. We replicate this in our controlled setup by targeting performance on the fixed set of discrete tasks in $\mathcal{D}_{\text{disc}}$, which requires the model to perform discrete regression (Equation 4). Fine-tuning is necessary since pretraining is both inefficient and limited by the distributional mismatch.
Figure 4: Fine-tuning hurts continuous loss. We train an $\alpha=0.2$ transformer with $64$ discrete tasks for $5000$ steps and fine-tune for $400$ steps on $\mathcal{D}_{\text{disc}}$ (highlighted). The discrete loss rapidly decreases, while the continuous loss rapidly increases.
@@
Fine-tuning the pretrained models from Section 2.4 over $\mathcal{D}_{\text{disc}}$ rapidly improves performance on $\mathcal{D}_{\text{disc}}$. However, this also leads to large performance drops on $\mathcal{D}_{\text{cont}}$ (pictured for $\alpha=0.5$ in Figure 4). We note that this forgetting is unnecessary since we can find solutions which do not exhibit this drastic increase (Appendix C.4). This is an instance of “catastrophic forgetting”, where fine-tuning a model to improve at one task causes it to worsen at other tasks. Though this performance drop could imply that the model can not perform ridge regression anymore, we investigate how fine-tuning is affecting model predictions to recover lost performance on the continuous distribution.